OpenStudy (anonymous):

State the domain and range of this square root function:y=3(sqrtx+2)+1 graphic. Then describe the transformations that you would do to get to the parent function.

OpenStudy (mathmale):

There are serious problems in your expression " y = 3(sqrtx+2)+1." Exactly what do you mean by "(sqrtx+2)" ? Which quantity is the "argument" of the sqrt function? Please re-write your question.

OpenStudy (anonymous):

it is written exactly how i need it to be plus its the way its written on my paper

OpenStudy (anonymous):

y= 3(sqrt x +2) +1

OpenStudy (anonymous):

domain \[x \geq 0\]

OpenStudy (anonymous):

when x=0 y=3(0+2)+1=7 range \[\geq 7\]